# Adapting Large Language Models for Document-Level Machine Translation

### 12 Jan 2024 v1

@@ Abstract
Large language models (LLMs) have made significant strides in various natural language processing (NLP) tasks. Recent research shows that the moderately-sized LLMs often outperform their larger counterparts after task-specific fine-tuning. In this work, we delve into the process of adapting LLMs to specialize in document-level machine translation (DocMT) for a specific language pair. Firstly, we explore how prompt strategies affect downstream translation performance. Then, we conduct extensive experiments with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our findings indicate that in some cases, these specialized models even surpass GPT-4 in translation performance, while they still significantly suffer from the off-target translation issue in others, even if they are exclusively fine-tuned on bilingual parallel documents. Furthermore, we provide an in-depth analysis of these LLMs tailored for DocMT, exploring aspects such as translation errors, the scaling law of parallel documents, out-of-domain generalization, and the impact of zero-shot crosslingual transfer. The findings of this research not only shed light on the strengths and limitations of LLM-based DocMT models but also provide a foundation for future research in DocMT.
@@ 1 Introduction
Large language models (LLMs) demonstrate impressive proficiency in a wide range of applications (Ouyang et al., 2022; Wei et al., 2022a; Sanh et al., 2022; Chung et al., 2022; OpenAI, 2023; Anil et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023). However, in the realm of translation tasks, only few very large models, such as gpt-3.5-turbo and gpt-4-turbo, can match or surpass the performance of state-of-the-art supervised encoder-decoder models like NLLB (Costa-jussà et al., 2022), while they still under-perform in translating low-resource languages (Robinson et al., 2023; Jiao et al., 2023; Hendy et al., 2023). Consequently, a number of recent works attempt to bridge the gap between LLMs and supervised encoder-decoder models in translation tasks (Zhu et al., 2023; Yang et al., 2023; Zhang et al., 2023; Moslem et al., 2023; Xu et al., 2023; Kudugunta et al., 2023). Recently, research suggests that smaller, specialized models can outperform larger, general-purpose models in specific tasks. (Gunasekar et al., 2023; Luo et al., 2023; Azerbayev et al., 2023). Therefore, we explore adapting moderately-sized LLMs for document-level machine translation (DocMT) in this study.
@@
In this work, we conduct a comprehensive empirical analysis on moderately-sized LLMs ($7B$ parameters in size) across 18 translation tasks involving nine language pairs. Our approach involves fine-tuning three moderately-sized LLMs using two strategies: Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT). We compare these models with state-of-the-art translation models, assessing them in terms of $s$BLEU, $d$BLEU, and COMET. Our results not only confirm the superior translation capabilities of LLMs after fine-tuning but also reveal a significant issue: off-target translations. This issue, commonly observed in zero-shot multilingual NMT models (Zhang et al., 2020a; Chen et al., 2023a), occurs in our LLM-based DocNMT models in some specific languages, even when the models are fine-tuned exclusively on bilingual corpora. Additionally, we present an in-depth analysis of our LLM-based DocNMT models from various perspectives: translation error distribution, the scaling law of parallel documents, out-of-domain generalization, and zero-shot cross-lingual transfer. This comprehensive study aims to advance understanding and improve the efficacy of LLMs in DocMT tasks.
@@
We present extensive empirical evidence that highlights both the superior translation capabilities and limitations of the LLM-based DocMT models in this study, making several significant discoveries that are crucial for future research. Here are the main takeaways:
  • Selective Excellence in Translation Tasks: Our empirical evidence reveals that our moderately-sized LLMs excel, even outperforming gpt-4-turbo, in specific translation tasks. However, they completely fail in others (Section 5). We find that the primary cause of these translation failures is the off-target translation problem (Section 6.1). When achieving similar performance, the LLM-based DocMT models produce fewer errors (Section 6.2).
@@
  • Fine-Tuning Strategies: We find that the PEFT approach yields superior overall performance compared to the FFT approach (Section 5). However, the FFT approach demonstrates better data efficiency, requiring only about $1\%$ of the total dataset (roughly $2K$ examples) to match the performance of models trained on the entire training set. In contrast, the PEFT approach requires $10\%$ of the total dataset (roughly $20K$ examples) to achieve comparable results (Section 6.3).
@@
  • Evaluation on Recent Test Sets: We evaluate our models on recent test sets between English and German from WMT2023 (Koehn et al., 2023). Our empirical results show that, when the data leakage risks are mitigated, the LLM-based DocMT models generalize better on out-of-domain text, compared to the conventional DocMT models (Section 6.4).
@@
  • Advantage of Base LLMs for Task-Specific Supervised Fine-Tuning: Our study shows that base LLMs, when used as the backbone for task-specific supervised fine-tuning, perform better than instruction-tuned LLMs. They demonstrate more effective zero-shot cross-lingual transfer, making them a preferable choice (Section 6.5).
@@ 2 Related Work
In recent years, numerous approaches have been proposed for document-level machine translation (DocMT). The widely used baseline approach consists of simply concatenating a few adjacent sentences and feeding this as an input to the MT system, without modifying the system architecture in any way (Tiedemann and Scherrer, 2017; Bawden et al., 2018; Nguyen et al., 2021; Sun et al., 2022). Apart from the simple concatenation method, there exist other approaches to DocMT, including document embedding (Macé and Servan, 2019; Huo et al., 2020), multiple encoders (Wang et al., 2017; Bawden et al., 2018; Voita et al., 2018; Zhang et al., 2018), attention variations (Miculicich et al., 2018; Zhang et al., 2020b; Maruf et al., 2019; Wong et al., 2020; Wu et al., 2023a), and translation caches (Maruf and Haffari, 2018; Tu et al., 2018; Feng et al., 2022). Furthermore, Maruf et al. (2022) present a comprehensive survey of DocMT.
@@
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Chowdhery et al., 2022; Scao et al., 2022; Anil et al., 2023; Touvron et al., 2023a, b). Furthermore, recent research has shown that supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) can significantly enhance their performance when following general language instructions Weller et al. (2020); Mishra et al. (2022); Wang et al. (2022); Wei et al. (2022a); Ouyang et al. (2022); Chung et al. (2022); Muennighoff et al. (2022); Wu et al. (2023b); Shen et al. (2023); Li et al. (2023); Wu and Aji (2023). More recently, there is a growing body of work exploring the translation capabilities of LLMs (Lu et al., 2023; Zhang et al., 2023; Xu et al., 2023; Robinson et al., 2023). However, it is important to note that these efforts have primarily focused on sentence-level machine translation (SenMT) and have not delved into document-level machine translation (DocMT). A noteworthy study in DocMT is conducted by Wang et al. (2023b), where they investigate the document-level translation capabilities of gpt-3.5-turbo, making it the most closely related work to our work.
@@ 3 Experimental Setup
In this study, we aim to adapt multilingual pre-trained large language models (LLMs) into a bilingual document-level machine translation (DocMT) model. In this section, we describe our experimental setup of this work, including training strategy (Section 3.1), datasets (Section 3.2), models (Section 3.3), evaluation methods (Section 3.4), and optimization (Section 3.5).
@@ 3.1 Two-Stage Training
DocMT approaches typically begin by pre-training the translation model on sentence-level parallel corpora, subsequently refining it through fine-tuning on document-level parallel corpora (Voita et al., 2019; Maruf et al., 2019; Ma et al., 2020; Sun et al., 2022; Wu et al., 2023a). More recently, Xu et al. (2023) propose a two-stage training strategy, which initially involves fine-tuning a LLM on monolingual text, followed by a second fine-tuning phase on parallel text. Given that most state-of-the-art open-sourced LLMs are trained on English-centric corpora, our approach begins with the fine-tuning of a LLM on monolingual documents, followed by fine-tuning on parallel documents. Following Xu et al. (2023), we omit the step of fine-tuning on sentence-level parallel datasets.
@@
Existing LLMs are typically pre-trained on English-centric corpora. Recent research highlights that these LLMs often exhibit sub-optimal performance on multilingual benchmarks (Li et al., 2023; Chen et al., 2023b; Scao et al., 2022). To address this limitation, our initial step involves fine-tuning LLMs using monolingual data from the target languages for each given language pair.
@@ 3.2 Datasets
Following Zhang et al. (2022), we conduct experiments on IWSLT2017 translation tasks (Cettolo et al., 2017). IWSLT2017 comprises translation datasets sourced from TED talks, encompassing translations between English and nine languages, including Arabic, German, French, Italian, Japanese, Korean, Dutch, Romanian, and Chinese. There are approximately $1.9K$ sentence-aligned parallel documents with about $240K$ sentences for each language pair. The dataset statistics can be found in Appendix A.
@@
Figure 1: Prompt types used in the preliminary study. <src_lang> and <tgt_lang> indicate the source and target languages. <src*> and <tgt*> indicate the source and target sentences. Note that the target sentences <tgt*> are only used during training and are replaced with the hypotheses <hyp*> generated by the model during inference. Concrete examples for each prompt variation can be found in Appendix B.
@@ 3.3 Models
  • Our re-implemented DocMT models: We conduct full fine-tuning on the concatenation-based DocMT model (Tiedemann and Scherrer, 2017), initialized with mT5 (Xue et al., 2021). These models are available with parameters of 300M, 580M, and 1.2B, representing the strong DocMT baseline.
@@
In this work, we utilize Llama2-7B, Bloom-7B, and Vicuna-7B, as our backbones. [^3] Llama2 signature: meta-llama/Llama-2-7b-hf, Bloom signature: bigscience/bloom-7b1, and Vicuna signature: lmsys/vicuna-7b-v1.5. Note that Vicuna-v1.5 models are fine-tuned from Llama2. The Llama2 models are predominantly pre-trained on English text, while the Bloom models are pre-trained on multilingual text. The use of Vicuna models allows us to compare the differences between base models and instruction-tuned models (Llama2 vs. Vicuna). We firstly fine-tune all the model parameters of these models on the collected monolingual documents, and then fine-tune these models on the parallel documents. We denote those fully fine-tuned models as L-7B-fft, B-7B-fft, and V-7B-fft. We denote those models fine-tuned with LoRA (Hu et al., 2022) as L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA.
@@ 3.4 Evaluation
Furthermore, as pointed out by Freitag et al. (2022) that conventional MT metrics like BLEU demonstrate poor correlation to human judgments, we also evaluate the translation quality with the state-of-the-art neural evaluation metric COMET Rei et al. (2020). [^5] COMET signature: Unbabel/wmt22-comet-da. Moreover, we use the average sentence-level BLEU $\mu_{s\textsc{BLEU}}$, the average document-level BLEU $\mu_{d\textsc{BLEU}}$, and the average COMET $\mu_{\textsc{COMET}}$ for the overall performance.
@@
3.5 Optimization and Hyperparameters
Fine-tuning on Monolingual Documents
Fine-tuning on Parallel Documents
@@ 3.5 Optimization and Hyperparameters
Table 1: Overall performance given by L-7B-LoRA and V-7B-LoRA on different prompt variations, across four English-centric translation tasks involving German and Chinese. PID indicates the prompt ID in Figure 1. Best results in each group are highlighted in bold.
@@ 4 A Preliminary Study on Prompts
Prompt is of vital importance in the research of LLMs. Recent studies demonstrate that an ideal prompt can significantly boost the model performance and unlock surprising model capabilities (Kojima et al., 2022; Wei et al., 2022b). Therefore, we firstly investigate impact of the prompt in the process of fine-tuning.
@@
As shown in Figure 1, we design four types of prompts in our preliminary study. With different designs of prompts, we aim to address two research questions: How does the context structure affect the translation quality? (Prompt 1 vs. Prompt 2) and How do the natural language instructions affect the translation quality? (Prompt 1 vs. Prompt 3). Furthermore, we explore the compound effect of these two aspects of our prompt designs in Prompt 4 (0(d)).
Results
Our investigation focuses on the impacts of prompt variations using three PEFT models: L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA, across four English-centric translation tasks involving German and Chinese. We present the overall results in Table 1. A comparison between Prompt 1 (0(a)) and Prompt 2 (0(b)) reveals that language models fine-tuned with Prompt 2 generally outperform those fine-tuned with Prompt 1. This suggests that Prompt 2’s structure, which incorporates extra contextual information in a few-shot-like context, is more effective for enhancing LLM performance. Regarding our second research question (0(a) vs. 0(c)), we observe a varied performance. Specifically, L-7B-LoRA and B-7B-LoRA exhibit better performance with Prompt 3 compared to Prompt 1. However, for V-7B-LoRA, we find the opposite trend: it performs better with Prompt 1 than with Prompt 3. These results demonstrate that the impact of prompt variations varies notably across different models. Furthermore, they suggest that natural language instructions are less effective when using instruction-tuned language models as model backbones. Finally, we observe that LLMs with Prompt 4 achieves the best overall performance, suggesting the positive compound effect of context structure and natural language instructions.
# of param. # of train. param. En-X X-En
$\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$ $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$
@@ Results
Bilingual LLM-based DocMT models (Ours)

### 15 Feb 2024 v2

@@ Abstract
Large language models (LLMs) have made significant strides in various natural language processing (NLP) tasks. Recent research shows that the moderately-sized LLMs often outperform their larger counterparts after task-specific fine-tuning. In this work, we delve into the process of adapting LLMs to specialize in document-level machine translation (DocMT) for a specific language pair. Firstly, we explore how prompt strategies affect downstream translation performance. Then, we conduct extensive experiments with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our findings indicate that in some cases, these specialized models even surpass GPT-4 in translation performance, while they still significantly suffer from the off-target translation issue in others, even if they are exclusively fine-tuned on bilingual parallel documents. Furthermore, we provide an in-depth analysis of these LLMs tailored for DocMT, exploring aspects such as translation errors, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluation on recent test sets, and zero-shot crosslingual transfer. Our findings not only shed light on the strengths and limitations of LLM-based DocMT models but also provide a foundation for future research.
@@ 1 Introduction
Large language models (LLMs) demonstrate impressive proficiency in a wide range of applications (Ouyang et al., 2022; Wei et al., 2022a; Sanh et al., 2022; Chung et al., 2022; OpenAI, 2023; Anil et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023). However, in the realm of translation tasks, only few very large models, such as gpt-3.5-turbo and gpt-4-turbo, can match or surpass the performance of state-of-the-art supervised encoder-decoder models like NLLB (Costa-jussà et al., 2022), while they still under-perform in translating low-resource languages (Robinson et al., 2023; Jiao et al., 2023; Hendy et al., 2023). Consequently, a number of recent works attempt to bridge the gap between LLMs and supervised encoder-decoder models in translation tasks (Zhu et al., 2023; Yang et al., 2023; Zhang et al., 2023; Moslem et al., 2023; Xu et al., 2023; Kudugunta et al., 2023). Recently, research suggests that smaller, specialized models can outperform larger, general-purpose models in specific tasks (Gunasekar et al., 2023; Luo et al., 2023; Azerbayev et al., 2023). Therefore, we explore adapting LLMs for document-level machine translation (DocMT) in this study.
@@
In this study, we analyze moderately-sized LLMs (with $7B$ parameters) across 18 translation tasks involving nine language pairs. We fine-tune three LLMs using Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT). Comparisons with state-of-the-art translation models, using metrics like $s$BLEU, $d$BLEU, and COMET, confirm the superior translation capabilities of LLMs after fine-tuning. However, we identify a significant issue of off-target translations, observed even after exclusive fine-tuning on bilingual corpora. Additionally, we present an in-depth analysis of our LLM-based DocNMT models from various perspectives: translation error distribution, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluations on WMT2023 test sets, and zero-shot cross-lingual transfer, aiming to enhance understanding and efficacy of LLMs in DocMT tasks.
@@
We present extensive empirical evidence that highlights both the superior translation capabilities and limitations of the LLM-based DocMT models in this study, making several significant discoveries. Here are the main takeaways:
  • Selective Excellence in Translation Tasks: Our findings show that our moderately-sized LLMs outperform gpt-4-turbo in certain translation tasks, but struggle in others due to the off-target translation issue. Despite this, our DocMT models exhibit better context awareness and fewer errors, while maintaining comparable performance.
@@
  • Fine-Tuning Strategies: Our research indicates that the PEFT approach outperforms the FFT approach overall. However, the FFT approach shows greater data efficiency, needing only about $1\%$ of the total dataset to reach the performance level of models trained on the entire dataset. In contrast, the PEFT approach requires $10\%$ of the total dataset for comparable results.
@@
  • Evaluation on Recent Test Sets: We evaluate our models on recent test sets between English and German from WMT2023 (Koehn et al., 2023). Our empirical results show that, when the data leakage risks are mitigated, the LLM-based DocMT models generalize better on out-of-domain text, compared to the conventional DocMT models.
@@
  • Advantage of Base LLMs for Task-Specific Supervised Fine-Tuning: Our study shows that base LLMs, when used as the backbone for task-specific supervised fine-tuning, perform better than instruction-tuned LLMs. They demonstrate more effective zero-shot cross-lingual transfer.
@@ 2 Related Work
In recent years, numerous approaches have been proposed for document-level machine translation (DocMT). There exist other approaches to DocMT, including document embedding (Macé and Servan, 2019; Huo et al., 2020), multiple encoders (Wang et al., 2017; Bawden et al., 2018; Voita et al., 2018; Zhang et al., 2018), attention variations (Miculicich et al., 2018; Zhang et al., 2020; Maruf et al., 2019; Wong et al., 2020; Wu et al., 2023), and translation caches (Maruf and Haffari, 2018; Tu et al., 2018; Feng et al., 2022). Furthermore, Maruf et al. (2022) present a comprehensive survey of DocMT.
@@
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Chowdhery et al., 2022; Scao et al., 2022; Anil et al., 2023; Touvron et al., 2023a, b). Furthermore, recent research has shown that supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) can significantly enhance their performance when following general language instructions Weller et al. (2020); Mishra et al. (2022); Wang et al. (2022); Shen et al. (2023); Li et al. (2023); Wu and Aji (2023). More recently, there is a growing body of work exploring the translation capabilities of LLMs (Lu et al., 2023; Zhang et al., 2023; Xu et al., 2023; Robinson et al., 2023). However, it is important to note that these efforts have primarily focused on sentence-level machine translation (SenMT) and have not delved into document-level machine translation (DocMT). A noteworthy study in DocMT is conducted by Wang et al. (2023b), where they investigate the document-level translation capabilities of gpt-3.5-turbo, making it the most closely related work to our work.
@@ 3 Experimental Setup
In this study, we aim to adapt multilingual pre-trained large language models (LLMs) into a bilingual document-level machine translation (DocMT) model. In this section, we describe our experimental setup of this work, including training strategy (Section 3.1), datasets (Section 3.2), models (Section 3.3), and evaluation (Section 3.4).
@@ 3.1 Two-Stage Training
DocMT approaches typically begin by pre-training the translation model on sentence-level parallel corpora, subsequently refining it through fine-tuning on document-level parallel corpora (Voita et al., 2019; Maruf et al., 2019; Ma et al., 2020; Sun et al., 2022; Wu et al., 2023). More recently, Xu et al. (2023) propose a two-stage training strategy, which initially involves fine-tuning a LLM on monolingual text, followed by a second fine-tuning phase on parallel text. Given that most state-of-the-art open-sourced LLMs are trained on English-centric corpora, our approach begins with the fine-tuning of a LLM on monolingual documents, followed by fine-tuning on parallel documents. Following Xu et al. (2023), we omit the step of fine-tuning on sentence-level parallel datasets.
@@
Existing LLMs are typically pre-trained on English-centric corpora. Recent research highlights that these LLMs often exhibit sub-optimal performance on multilingual benchmarks (Li et al., 2023; Chen et al., 2023; Scao et al., 2022). To address this limitation, our initial step involves fine-tuning all the parameters of LLMs using monolingual data from the target languages.
@@ 3.2 Datasets
Following Zhang et al. (2022), we conduct experiments on IWSLT2017 translation tasks (Cettolo et al., 2017). IWSLT2017 comprises translation datasets sourced from TED talks, encompassing translations between English and nine other languages, including Arabic, German, French, Italian, Japanese, Korean, Dutch, Romanian, and Chinese. There are approximately $1.9K$ sentence-aligned parallel documents with about $240K$ sentences for each language pair. The dataset statistics can be found in Appendix A.
@@
Figure 1: Prompt types used in the preliminary study. <src_lang> and <tgt_lang> indicate the source and target languages. <src*> and <tgt*> indicate the source and target sentences. Note that the target sentences <tgt*> are only used during training and are replaced with the hypotheses <hyp*> generated by the model during inference. Concrete examples for each prompt variation can be found in Appendix C.
@@ 3.3 Models
  • Our re-implemented DocMT models: We conduct full fine-tuning on the concatenation-based DocMT model (Tiedemann and Scherrer, 2017), as well as several recent DocMT baselines (Sun et al., 2022; Wu et al., 2023, 2024), initialized with mT5 (Xue et al., 2021). These models are available with parameters of 300M, 580M, and 1.2B, representing the strong DocMT baseline.
@@
In this work, we utilize Llama2-7B, Bloom-7B, and Vicuna-7B, as our backbones. [^3] Llama2 signature: meta-llama/Llama-2-7b-hf, Bloom signature: bigscience/bloom-7b1, and Vicuna signature: lmsys/vicuna-7b-v1.5. Note that Vicuna-v1.5 models are fine-tuned from Llama2. The Llama2 models are predominantly pre-trained on English text, while the Bloom models are pre-trained on multilingual text. The use of Vicuna models allows us to compare the differences between base models and instruction-tuned models (Llama2 vs. Vicuna). We denote those fully fine-tuned models as L-7B-fft, B-7B-fft, and V-7B-fft. We denote those models fine-tuned with LoRA (Hu et al., 2022) as L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA. The optimization details can be found in Appendix B.
@@ 3.4 Evaluation
Furthermore, as conventional MT metrics like BLEU demonstrate poor correlation to human judgments (Freitag et al., 2022), we also evaluate the translation quality with the state-of-the-art neural evaluation metric COMET Rei et al. (2020). [^5] COMET signature: Unbabel/wmt22-comet-da. Moreover, we use the average sentence-level BLEU $\mu_{s\textsc{BLEU}}$, the average document-level BLEU $\mu_{d\textsc{BLEU}}$, and the average COMET $\mu_{\textsc{COMET}}$ for the overall performance.
@@
Table 1: Overall performance given by L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA on different prompt variations, across four English-centric translation tasks involving German and Chinese. PID indicates the prompt ID in Figure 1. Best results are highlighted in bold.
@@ 4 A Preliminary Study on Prompts
The prompt plays a crucial role in LLM research. Recent studies show that an optimal prompt can greatly enhance model performance and reveal unexpected model capabilities (Kojima et al., 2022; Wei et al., 2022b). Hence, our initial focus is on investigating the prompt’s impact during fine-tuning.
@@
Displayed in Figure 1, our preliminary study features four prompt types. These designs aim to tackle two research questions: How does context structure impact translation quality? (Prompt 1 vs. Prompt 2) and How do natural language instructions influence translation quality? (Prompt 1 vs. Prompt 3). We also investigate the combined effect of these aspects in Prompt 4.
# of param. # of train. En-X X-En
param. $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$ $\mu_{s\textsc{BLEU}}$ $\mu_{d\textsc{BLEU}}$ $\mu_{\textsc{COMET}}$
@@
MR-Doc2Sen-mT5 (2022) 01.2B 01.2B 18.8 21.9 79.9 22.0 23.8 79.3
MR-Doc2Doc-mT5 (2022) 01.2B 01.2B 22.5 24.0
DocFlat-mT5 (2023) 01.2B 01.2B 19.2 22.4 80.2 22.2 24.3 79.3
IADA-mT5 (2024) 01.2B 01.2B 19.3 22.4 80.4 22.1 24.0 79.5
LLM-based DocMT models (Ours)
@@
Results
Our investigation analyzes prompt variations using three PEFT models (L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA) on four English-centric translation tasks involving German and Chinese. Overall results are presented in Table 1. Comparing Prompt 1 (0(a)) and Prompt 2 (0(b)), we find that models fine-tuned with Prompt 2 generally outperform those with Prompt 1, indicating Prompt 2’s effectiveness in enhancing LLM performance. Regarding our second research question (0(a) vs. 0(c)), we observe varied performance. L-7B-LoRA and B-7B-LoRA perform better with Prompt 3, while V-7B-LoRA performs better with Prompt 1. These results highlight varying impacts of prompt variations across models and suggest natural language instructions are less effective when using instruction-tuned language models as model backbones. Finally, LLMs with Prompt 4 (0(d)) achieve the best overall performance, suggesting a positive compound effect of context structure and instructions.